Grade upto college level General Physics

The navigator of an oil tanker uses the satellites of the Global Positioning System (GPS/NAVSTAR) to find latitude and longitude; see Fig. These are 43036’25.3” N and 77031’48.2” W. If the accuracy of these determinations is ±0.5”, what is the uncertainty in the tanker’s position measured along (a) a north-south line (meridian of longitude) and (b) an east-west line (parallel of latitude)? (c) Where is the tanker?

Amit Saxena
11 Years agoGrade upto college level

1 Answer

Navjyot Kalra
11 Years ago
(a)
The circumference S of the Earth is approximately
S = 40, 000 km
Now, \pm0.5’‘ of an arc with respect to the circumference of the Earth is
\theta0 = \frac{0.5}{(60\times 60\times 360)}
= 3.9 x 10-7
Substitute 40,000 for S in the equation \Delta E = \pm \theta0.5S
\Delta E = \pm \theta0.5S
= \pm(3.9\times 10^{-7}) (40,000 km) (\frac{10^{3} m}{1km} )
=\pm 15.6m

Then, the range of the error is
R_1{= 2 \Delta E}
= 2 (\pm 15.6m)
= \pm 31.2 m

Rounding off to two significant figures, the uncertainty in the tanker’s position measured along the north-south line is . 31 m.
(b)
The east-west range is smaller because the distance measured along latitude is smaller than the circumference by a factor of the cosine of the latitude.
Substitute 31m for R1 and 43.60 for\theta in the equation . R = R1 cos\theta.
R = R1 cos \theta
= (31 m) (cos 43.60)
= 22.449 m
Rounding off to two significant figures, the uncertainty in the tanker’s position measured along the east-west line is 22 m.
(c)
The tanker is in Lake Ontario. It is about 20 km off the coast of Hamlin.